RC meeting and Run Plan, Wednesday Sept. 13, 2023
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
- CG: always call the RC first
- Andrew L: for the Hall A access procedure needs to filled out before you do the access. There are copies in the folder downstairs.
- Mark: we should scan them an put them in the log (picture with phone is ok)
- AL: changes to the procedure: any work inside the SBS, BB boundary need to have the PS locked out. Two QEWs need to be present to throw the big switch. With the exception of field measurements.
- AL: hall A call-in instructions. Techs on site 7-4:30 Mon-Th, 7-11AM Fr -> during these times call Zak first.
- AL: everything goes through the run coordinator (The techs will not answer unknown numbers).
- Mark: call the RC all the time!
- AL: after 9pm everything non emergency needs to be deferred. Call RC, who will call Mark and they will decide
- Operational restrictions: BH: probably no changes needed
Readiness for Beam
- CG: Working through the startup procedure. Completed a series of carbon hole measurements and an NMR measurements (21% polarization for 5 hours worth of pumping). Plan is to take a quick escorted access to recovery layer 0 of BB GEM HV and the BBCal
- What are the decision thresholds for laser work? GC: below 50-52% polarization we should go down. However, overnight once we get 30-35% we can start production. 2H in CANS mode -> no other work in the hall.
- When do we want to have a Moller polarization measurement? BW: decision after target polarization is established
- Outstanding issues:
- laser polarization checks - keep a controlled access as a possibility for tomorrow
- Moller quads: there is a ground leak from the LCW system. Plan is to let them flush (they are NOT needed for optics, but needed for a Moller measurement) and we will check on them during an opportunistic down or when we absolutely need to have a Moller measurement
- Moller target magnet: needs to be trained for ~8h before a measurement can be done. This prevents beam delivery so we need to be careful about scheduling this.
- SBS Left rack: breaker replaced and the setup was looked over by Joe Beaufait. At an opportune time we need facilities to fix the type of breaker they are using for our two phase supply. Alex and Ellen evened out the power draw among the two phases.
- CG: shifts populated. Faraz will take tomorrow swing if no MOLLER measurement on Friday.
For each subsystem we ask that the responsible person go over the following points:
- The overall status of your subsystem
- Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
- Does your system being ready depend on your team only or are outside resources required?
- SBS magnet and Bigbite magnets
- BigBite magnet cannot be turned off or on remotely (https://logbooks.jlab.org/entry/4178242). Ellen fixed it => confirm
- Corrector magnets:
- work as expected
- Target status (magnets, measurements, laser, installation, alignment, polarization, power etc)
- 21% polarized
- Hunter did the EPR calibration and working on better understanding. Rough number for future NMR calibrations is 1.1
- HCal (Jiwan Poudel)
- VME DAC GUI does not read the correct threshold; confirm with Alex about fix
- Bigbite (Arun Tadepalli)
- we have the correct values loaded now (see recent log entry from this morning)
- Counting House Alarms (Bill Henry)
- Bill tested the alarms and instructions for shift crew response in the wiki (needs further revision?)
- DAQ (Alex Camsonne)
- GEMs (Holly Szumila-Vance)
- ChG: looked at online plots and looked fine (experts should confirm). AP: BB GEMs looked ok
- ChG: Ellen has a mass flow controller and needs to find some time to install it (not a show stopper - gas mixture changes from 79%Ar to 70%) => Chandan can also change the MFC; this is outside of the hall can be done later
- Moller polarimeter (Don Jones)
- Rachel: EPICS communications issues yesterday; the HV is on but one PMT is not reaching the full voltage (might want to take a look - not urgent since it's out of the acceptance).
Preliminary run plan
- First, bring the beam safely into the hall (Beam recovery procedure - Jiwan and William Henry https://sbs.jlab.org/wiki/index.php/HOW_TOs#Beam_Recovery_Procedure)
- Target performance assessment by target group - Arun (Why hasn’t the target reached xx% polarization?), tell us when we will get there (Target group - Arun) and plan for those hours in between
- Take production data!
- Production: 1h for 5, 10, 20, 30uA => run 30uA until target experts confirm everything is as it should be. These will be singles (not coincidence). PS5 should be set to 2 for helicity readback.
- Charge asymmetry: Paul will run feedback once a day until hall C comes up.
List of items in case target isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- Moller measurement depending on availability of people and results from Hall C
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)
In person: Ciprian G, Sanghwa P, Mark J, Chandan G, Andrew L
Online: Ole H, Alexandre C, Bogdan W, Sean J, Jimmy C, Paul K, Gordon C, Rachel M, Jiwan P, Silviu CD, Andrew P, Jack J,